# references.bib

### v1

@@ 5 Numerical experiments: A nonlinear system of two resonators
In Figure 2(b) and Figure 2(c), the solutions for asymmetric systems are computed. The domain $B_{2}$ is changed to $B_{0.21}(\frac{1}{2}e_{3})$ and $B_{0.22}(\frac{1}{2}e_{3})$, respectively. For $\lvert(q_{0})_{1}\rvert$, $\lvert(q_{0})_{2}\rvert$ small enough, one observes that the solution curves are close to the solutions of the linear setting, as one would expect from linearization techniques. Due to the asymmetry of the system, the solutions $q_{0}$ of the linear system have entries of different magnitudes $\lvert(q_{0})_{1}\rvert\not=\lvert(q_{0})_{2}\rvert$, which makes the two subspaces visible (light blue and red curves no longer lie on top of each other, as in Figure 1(a)). For sufficiently large magnitudes, a third solution appears. In the symmetric setting, the third solution did not interfere with the ‘linear’ solutions predicted by Lemma 5.1, when the symmetry is broken, the third solution interferes with the first two. This can be seen in the splitting of the light blue curve.
Figure 1: In these plots the solution curves of equation (20) for different choices of $\operatorname{{\mathcal{D}}}$ are plotted. From left to right, $\operatorname{{\mathcal{D}}}=B_{0.2}\left(-\frac{1}{2}e_{3}\right)\cup B_{0.2}\left(\frac{1}{2}e_{3}\right)$, $\operatorname{{\mathcal{D}}}=B_{0.2}\left(-\frac{1}{2}e_{3}\right)\cup B_{0.21}\left(\frac{1}{2}e_{3}\right)$ and $\operatorname{{\mathcal{D}}}=B_{0.2}\left(-\frac{1}{2}e_{3}\right)\cup B_{0.22}\left(\frac{1}{2}e_{3}\right)$, respectively. The remaining parameters are given by $\beta=0.1$ and ${c_{r}}=1$. The curves trace the magnitude of the entries of different solutions $q_{0}$ to equation (20). The colors account for the complex valued solutions $q_{0}\in\mathbb{C}^{2}$ and depict the phase of $(q_{0})_{1}/(q_{0})_{2}$.
@@
The respective subwavelength resonances of the nonlinear setting are displayed in Figure 2. The subwavelength resonances of the symmetric setting associated to Figure 1(a) are depicted in Figure 2(a). In the main plot the value of subwavelength resonances are plotted in the complex plane. In the inlet plot the subwavelength solution with the colors of the different solution curves corresponding to the different curves of the subwavelength resonances are plotted. In Figure 2(a), one observes that the curve of the subwavelength resonances which are associated to the third solution curve, touches the curve of the subwavelength resonances associated to the solution $q_{0}=(a,a)^{\top},a\in\mathbb{C}$, and the superscript $\top$ denotes the transpose. It is precisely that curve that is mixing with the nonlinear solutions for sufficiently large amplitudes in the asymmetric settings, as can be seen from Figures 1(b) and 2(b), and Figures 1(c) and 2(c), respectively.
@@
Figure 2: In this figure the subwavelength resonances associated to the settings discussed in the caption of Figure 1 are displayed. The resonances are displayed as curves in the complex plane. The color of each curve is associated to the color of a solution curve depicted in the inlet plots.
@@ 6 Conclusion and outlook
In this paper, we have provided for the first time a discrete approximation to the nonlinear resonance problem for finite systems of subwavelength resonators. Our discrete approximation allows not only deriving the effect of weak nonlinearities on the linear resonances but more importantly characterizes the soliton-like solutions which are induced by stronger nonlinearities. It would be very interesting, firstly, to rigorously derive modal decompositions for scattered solutions by nonlinear systems of resonators and, secondly, to analyze the stability of the soliton-like resonances and their associated eigenmodes. Our results in this paper can be extended to periodic systems of resonators. We plan in forthcoming works to elucidate the interplay between nonlinearities and subwavelength resonances, in particular the nonlinearity-induced topological transitions in such systems.
@@ Acknowledgments
This work was supported in part by the Swiss National Science Foundation grant number 200021–200307. T.K. would like to thank Tianwei Yu and Wouter Tonnon for the numerous discussions.
[…]

### v3

@@ 5 Numerical experiments: A nonlinear system of two resonators
In Figure 1(b) and Figure 1(c), the solutions for asymmetric systems are computed. The domain $B_{2}$ is changed to $B_{0.21}(\frac{1}{2}e_{3})$ and $B_{0.22}(\frac{1}{2}e_{3})$, respectively. For $\lvert(q_{0})_{1}\rvert$, $\lvert(q_{0})_{2}\rvert$ small enough, one observes that the solution curves are close to the solutions of the linear setting, as one would expect from linearization techniques. Due to the asymmetry of the system, the solutions $q_{0}$ of the linear system have entries of different magnitudes $\lvert(q_{0})_{1}\rvert\not=\lvert(q_{0})_{2}\rvert$, which makes the two subspaces visible (light blue and red curves no longer lie on top of each other, as in Figure 1(a)). For sufficiently large magnitudes, a third solution appears. In the symmetric setting, the third solution did not interfere with the ‘linear’ solutions predicted by Lemma 5.1. However, when the symmetry is broken, the third solution interferes with the first two. This can be seen in the splitting of the light blue curve.
@@
The leading order asymptotics of the respective subwavelength resonances of the nonlinear setting are displayed in Figure 5.2. The leading order asymptotics of the subwavelength resonances of the symmetric setting associated to Figure 1(a) are depicted in Figure 2(a). In the main plot, the value of the leading order asymptotics $\omega_{0}$ of the subwavelength resonances are plotted in the complex plane. In the inlet plot the leading order asymptotics $q_{0}$ of the subwavelength solution with the colors of the different solution curves corresponding to the different curves of the leading order asymptotics of the subwavelength resonances are plotted. In Figure 2(a), one observes that the curve of the leading order asymptotics of the subwavelength resonances which are associated to the third solution curve, touches the curve of the leading order asymptotics of the subwavelength resonances associated to the solution $q_{0}=(a,a)^{\top},a\in\mathbb{C}$, where the superscript $\top$ denotes the transpose. It is precisely that curve that is mixing with the nonlinear solutions for sufficiently large amplitudes in the asymmetric settings, as can be seen from Figures 1(b) and 2(b), and Figures 1(c) and 2(c), respectively.
@@
Figure 5.2: In this figure the leading order asymptotics of the subwavelength resonances associated to the settings discussed in the caption of Figure 5.1 are displayed. The leading order asymptotics of the resonances are displayed as curves in the complex plane. The color of each curve is associated to the color of a solution curve $q_{0}$ depicted in the inlet plots.
@@ 6 Conclusion and outlook
In this paper, we have provided for the first time a discrete approximation to the nonlinear resonance problem for finite systems of subwavelength resonators. Our discrete approximation allows us not only to derive the effect of weak nonlinearities on the linear resonances but more importantly characterizes the solutions which are induced by stronger nonlinearities.
It would be very interesting, firstly, to rigorously derive modal decompositions for scattered solutions by nonlinear systems of resonators, and secondly, to prove existence of soliton-like resonances and their associated localized eigenmodes and analyze their stability. Another very interesting problem is to develop a Foldy-Lax formulation or a point-interaction approximation for nonlinear subwavelength resonators and study their limits in large, dilute systems as done in the linear case, for instance, in [6, 2]. Our results in this paper can be extended to periodic systems of resonators. In forthcoming works, we plan on the one hand to prove existence of soliton-like solutions in periodic structures that are induced by nonlinear defects into a linear structure and on the other hand, to elucidate the interplay between nonlinearities and subwavelength resonances, in particular the nonlinearity-induced topological transitions in such systems.
@@ Acknowledgments
This work was supported in part by the Swiss National Science Foundation grant number 200021–200307. T.K. thanks Tianwei Yu, Wouter Tonnon and Michael Weinstein for the interesting discussions.